Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina ( at Mon Nov 10 04:57:38 MET 1997) Minor failures
some minor problems

Effelsberg (dgraham at Mon Nov 10 09:43:24 MET 1997) Failure
No useful data. Wrongly translated using drudg to use USB of VCs.

HartRAO (Jonathan Quick at Mon Nov 10 09:46:47 MET 1997) Minor failures
Missed one scan (50mins).

Jodrell (Alastair Gunn at Mon Nov 10 12:40:27 MET 1997) Minor failures
Hopefully the LOs were set to give the correct frequencies (we did our best). Track 21 dead. No other known problems.

Torun (K.Borkowski at Mon Nov 10 12:47:36 MET 1997) Success

Westerbork (Tony Foley at Mon Nov 10 14:13:33 MET 1997) Minor failures
Looks OK, but note that the frequency is at the bottom edge of the frontend passband. Only VC1 to VC4 have internal 500kHz filters and I connected external 500kHz filters to VC5 VC6 and VC7 USB only. Other VCs have no signal.

Noto (C.Trigilio at Fri Nov 14 09:34:36 MET 1997) Success

Calibration (Fredrik T Rantakyro at Tue Dec 9 14:56:15 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name cah1a.antab Read the comments in the file to get further information.